# Continuous-time Risk-sensitive Reinforcement Learning via Quadratic Variation Penalty

### v1

@@ 6 Conclusion
There are alternative formulations of risk sensitivity other than the exponential form proposed in discrete-time systems, e.g., Xu et al., (2023); Wu and Xu, (2023), however, it still seems unclear what the continuous-time counterpart is and if any form of penalty can be introduced to transform the nonlinear objective to a recursive utility maximization problem.
@@ Acknowledgement
The author is supported by the Start-up Fund and the Faculty Direct Fund 4055220 at The Chinese University of Hong Kong.
@@ A.2 Ergodic linear-quadratic control
Conjecturing $\varphi(x)=\frac{1}{2}k_{2}x^{2}+k_{1}x$ and plugging it to the HJB equation, we get the first-order condition $a^{*}=\frac{[k_{2}(B+CD)-R]x+k_{1}B-Q}{N-k_{2}D^{2}}$, assuming $N-k_{2}D^{2}>0$. The HJB equation now becomes
@@
This leads to three algebraic equations by matching the coefficients of $x^{2}$, $x$ and the constant term:
[…]
@@ B.1 Proof Lemma 1
Suppose $\{Z_{s},s\geq 0\}$ is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$- adapted continuous semimartingale with $\mathbb{E}[e^{\frac{\epsilon^{2}}{2}\langle Z\rangle(s)}]<\infty$ for every $s\in[0,T]$.
@@
    If $\{Z_{s}+\frac{\epsilon}{2}\langle Z\rangle(s),s\geq 0\}$ is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local martingale, then $\{e^{\epsilon Z_{s}},\geq 0\}$ is also a ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$ martingale.
@@
    If $\epsilon>0$, $\{Z_{s}+\frac{\epsilon}{2}\langle Z\rangle(s),s\geq 0\}$ is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local submartingale (supermartingale), then $\{e^{\epsilon Z_{s}},s\geq 0\}$ is also a ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$ submartingale (supermartingale).
@@
    If $\epsilon<0$, $\{Z_{s}+\frac{\epsilon}{2}\langle Z\rangle(s),s\geq 0\}$ is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local submartingale (supermartingale), then $\{e^{\epsilon Z_{s}},s\geq 0\}$ is also a ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$ supermartingale (submartingale).
@@ B.2 Proof of Theorem 1
is an ($(\{\mathcal{F}_{s}^{X}\}_{s\geq 0},\mathbb{P})$- (local) martingale, we have
@@
×$\times\boldsymbol{\pi}^{*}(a|t,x)\mathrm{d}a.$
@@ B.4 Proof of Theorem 2
    $\hat{r}(t,x,a):=$ $\frac{\partial\widehat{J^{*}}}{\partial t}(t,x)+b(t,x,a)\circ\frac{\partial\widehat{J^{*}}}{\partial x}(t,x)+\frac{1}{2}\sigma\sigma^{\top}(t,x,a)\circ\frac{\partial^{2}\widehat{J^{*}}}{\partial x^{2}}(t,x)$
@@
    is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local martingale, which follows from applying Itô’s lemma to the above process. As a result, $\int_{t}^{s}[r(u,{X}_{u}^{\boldsymbol{\pi}},a^{\boldsymbol{\pi}}_{u})-\widehat{q^{*}}(u,{X}_{u}^{\boldsymbol{\pi}},a^{\boldsymbol{\pi}}_{u})+\hat{r}(u,{X}_{u}^{\boldsymbol{\pi}},a^{\boldsymbol{\pi}}_{u})]\mathrm{d}u$ is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local martingale.
@@
    is an ($(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$- (local) martingale, which follows from applying Itô’s lemma to the above process. Hence it is also an ($(\{\mathcal{F}_{s}^{X}\}_{s\geq 0},\mathbb{P})$- (local) martingale because $\mathcal{F}_{s}^{X}$ is a sub-sigma field of $\mathcal{F}_{s}$.
@@ B.6 Proof of Theorem 4
Suppose there are sequences of non-negative numbers $\{e_{i}\}_{i\geq 0},\{a_{i}\}_{i\geq 0},\{\eta_{i}\}_{i\geq 0}$ such that $a_{i}\in(0,1)$, $e_{i+1}\leq(1-a_{i})e_{i}+a_{i}^{2}\eta_{i}$ for all $i\geq 0$. If for all $i\geq 0$, with $a_{i}\leq a_{i+1}(1+Aa_{i+1})$ for some constant $10<A<$ 1, satisfying $\frac{1}{1-A}\geq\frac{(1-a_{0})e_{0}}{a_{0}\eta_{0}}+a_{0}$, and $\eta_{i}$ is non-decreasing in $i$, then $e_{i+1}\leq\frac{1}{1-A}a_{i}\eta_{i}$ for all $i\geq 0$. In particular, for any $p\in(\frac{1}{2},1]$, a sequence in the form of $a_{i}=\frac{\alpha}{(i+\beta)^{p}}$ with suitable choice of $\alpha,\beta$ satisfies $a_{i}\leq a_{i+1}(1+Aa_{i+1})$ for all $i\geq 0$.
@@ B.6.1 Stage 1: Properties of the increment
$=$ $\left[r+(\mu-r)a(t)-\frac{\gamma}{2}a(t)^{2}\sigma^{2}\right]\mathrm{d}t+a_{i}(t)\sigma\mathrm{d}W(t)-\theta\mathrm{d}t$

### v2

@@ 6 Discussions
There are alternative formulations of risk sensitivity other than the exponential form proposed in discrete-time systems, e.g., Xu et al., (2023); Wu and Xu, (2023), however, it still seems unclear what the continuous-time counterpart is and if any form of penalty can be introduced to transform the nonlinear objective to a recursive utility maximization problem. How to address the distributional robustness other than the KL divergence in the continuous-time diffusion processes in a tractable way and to conduct RL still remains largely an open question for future research.
@@ Acknowledgement
The author is supported by the Start-up Fund at The Chinese University of Hong Kong and Hong Kong Research Grants Council (RGC) - Early Career Scheme (ECS) 24211124. I thank the participants at the ETH-Hong Kong-Imperial Mathematical Finance Workshop, the 1st INFORMS Conference on Financial Engineering and FinTech, and 2025 SIAM Conference on Financial Mathematics and Engineering for their helpful discussions. I also thank two anonymous reviewers for their constructive and detailed comments that have led to an improved version of this paper.
@@ A.2 Ergodic linear-quadratic control
Conjecturing $\varphi(x)=\frac{1}{2}k_{2}x^{2}+k_{1}x$ and plugging it into the HJB equation, we get the first-order condition $a^{*}=\frac{[k_{2}(B+CD)-R]x+k_{1}B-Q}{N-k_{2}D^{2}}$, assuming $N-k_{2}D^{2}>0$. The HJB equation now becomes
@@
This leads to three algebraic equations by matching the coefficients of $^{2}, xx$ and the constant term:
$\left\{\begin{aligned} &k_{2}(2A+C^{2})-M+\frac{[k_{2}(B+CD)-R]^{2}}{N-k_{2}D^{2}}=0,\\ &k_{1}A-P+\frac{[k_{2}(B+CD)-R](k_{1}B-Q)}{N-k_{2}D^{2}}=0,\\ &\beta=\frac{(k_{1}B-Q)^{2}}{2(N-k_{2}D^{2})}.\end{aligned}\right$.
@@ B.1 Proof Lemma 1
Suppose $\{Z_{s},s\geq 0\}$ is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$- adapted continuous semimartingale with $\mathbb{E}[e^{\frac{\epsilon^{2}}{2}\langle Z\rangle(s)}]<\infty$ for every $s\in[0,T]$.
@@
    If $\{Z_{s}+\frac{\epsilon}{2}\langle Z\rangle(s),s\geq 0\}$ is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local martingale, then $\{e^{\epsilon Z_{s}},\geq 0\}$ is also an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$ martingale.
@@
    If $\epsilon>0$, $\{Z_{s}+\frac{\epsilon}{2}\langle Z\rangle(s),s\geq 0\}$ is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local submartingale (supermartingale), then $\{e^{\epsilon Z_{s}},s\geq 0\}$ is also an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$ submartingale (supermartingale).
@@
    If $\epsilon<0$, $\{Z_{s}+\frac{\epsilon}{2}\langle Z\rangle(s),s\geq 0\}$ is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local submartingale (supermartingale), then $\{e^{\epsilon Z_{s}},s\geq 0\}$ is also an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$ supermartingale (submartingale).
@@ B.2 Proof of Theorem 1
is an $(\{\mathcal{F}_{s}^{X}\}_{s\geq 0},\mathbb{P})$- (local) martingale, we have
@@
$\times\boldsymbol{\pi}^{*}(a|t,x)\mathrm{d}a.$
@@ B.4 Proof of Theorem 2
    $\hat{r}(t,x,a)$= $\frac{\partial\widehat{J^{*}}}{\partial t}(t,x)+b(t,x,a)\circ\frac{\partial\widehat{J^{*}}}{\partial x}(t,x)+\frac{1}{2}\sigma\sigma^{\top}(t,x,a)\circ\frac{\partial^{2}\widehat{J^{*}}}{\partial x^{2}}(t,x)$
@@
    is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local martingale, which follows from applying Itô’s lemma to the above process. As a result, $\int_{t}^{s}[r(u,{X}_{u}^{\boldsymbol{\pi}},a^{\boldsymbol{\pi}}_{u})-\widehat{q^{*}}(u,{X}_{u}^{\boldsymbol{\pi}},a^{\boldsymbol{\pi}}_{u})+\hat{r}(u,{X}_{u}^{\boldsymbol{\pi}},a^{\boldsymbol{\pi}}_{u})]\mathrm{d}u$ is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$-local martingale.
@@
    is an $(\{\mathcal{F}_{s}\}_{s\geq 0},\mathbb{P})$- (local) martingale, which follows from applying Itô’s lemma to the above process. Hence it is also an $(\{\mathcal{F}_{s}^{X}\}_{s\geq 0},\mathbb{P})$- (local) martingale because $\mathcal{F}_{s}^{X}$ is a sub-sigma field of $\mathcal{F}_{s}$.
@@ B.6 Proof of Theorem 4
Suppose there are sequences of non-negative numbers $\{e_{i}\}_{i\geq 0},\{a_{i}\}_{i\geq 0},\{\eta_{i}\}_{i\geq 0}$ such that $a_{i}\in(0,1)$, $e_{i+1}\leq(1-a_{i})e_{i}+a_{i}^{2}\eta_{i}$ for all $i\geq 0$. If for all $i\geq 0$, with $a_{i}\leq a_{i+1}(1+Aa_{i+1})$ for some constant $0<A<1$, satisfying $\frac{1}{1-A}\geq\frac{(1-a_{0})e_{0}}{a_{0}\eta_{0}}+a_{0}$, and $\eta_{i}$ is non-decreasing in $i$, then $e_{i+1}\leq\frac{1}{1-A}a_{i}\eta_{i}$ for all $i\geq 0$. In particular, for any $p\in(\frac{1}{2},1]$, a sequence in the form of $a_{i}=\frac{\alpha}{(i+\beta)^{p}}$ with suitable choice of $\alpha,\beta$ satisfies $a_{i}\leq a_{i+1}(1+Aa_{i+1})$ for all $i\geq 0$.
@@
To get the big picture of how the proof is conducted, I first illustrate the ideas by assuming ideal continuous sampling and computation of integral are feasible in the following Stages 1 and 2. Then I give the proof by considering the discretization error in Stage 3.
@@ B.6.1 Stage 1: Properties of the increment
$=$ $\left[r+(\mu-r)a(t)-\frac{\gamma}{2}a(t)^{2}\sigma^{2}\right]\mathrm{d}t+a(t)\sigma\mathrm{d}W(t)-\theta\mathrm{d}t$
@@ B.6.2 Stage 2: Recursive relations
B.6.3 Stage 3: Impact of Discretization
In the above analysis, I ignored the error caused by discretization. Here, the errors caused by time discretization are twofold. First, continuously sampling i.i.d. actions $a(t)$ is practically infeasible, and also has measure-theoretical issues (see Remark 2.1 in Szpruch et al., (2024) for more details). Second, the integral can only be approximated by suitable finite sums in numerical computation.
Therefore, let us consider the dynamics of the wealth equation with discretized time grid: $0=t_{0}<t_{1}<\cdots<t_{K}=T$ such that for $t\in[t_{k-1},t_{k}]$ with $k=1,\cdots,K$,
$\mathrm{d}X(t)=[r+(\mu-r)a(t_{k-1})]X(t)\mathrm{d}t+X(t)a(t_{k-1})\sigma\mathrm{d}W(t).$
For simplicity, here, the time grid is equally spaced, i.e., $\Delta t=\frac{T}{K}$ and $t_{k}=k\Delta t$. Note that this is the same wealth equation as (20) with a piecewise constant action sequence $a(t_{0}),a(t_{1}),\cdots,a(t_{K-1})$. On the interval $t_{k-1},t_{k}]$, we can still apply Itô’s lemma to the TD error (34), i.e.,
$V\left(t_{k},X(t_{k});\theta\right)-V\left(t_{k-1},X(t_{k-1});\theta\right)-q\left(a(t_{k-1});\boldsymbol{\psi}\right)\Delta t$
$+\frac{1-\gamma}{2}\left(V\left(t_{k},X(t_{k});\theta\right)-V\left(t_{k-1},X(t_{k-1});\theta\right)\right)^{2}$
$=$ $-\theta\Delta t+\log X(t_{k})-\log X(t_{k-1})+\left[\frac{(a(t_{k-1})-\psi_{1})^{2}}{2\psi_{2}}+\frac{\lambda}{2}\log 2\pi\lambda+\frac{\lambda}{2}\log\psi_{2}\right]\Delta t$
$+\frac{1-\gamma}{2}\left(-\theta\Delta t+\log X(t_{k})-\log X(t_{k-1})\right)^{2}$
$=$ $\left[-\theta+r+(\mu-r)a(t_{k-1})-\frac{1}{2}\sigma^{2}a(t_{k-1})^{2}\right]\Delta t+\sigma a(t_{k-1})\left[W(t_{k})-W(t_{k-1})\right]$
$+\left[\frac{(a(t_{k-1})-\psi_{1})^{2}}{2\psi_{2}}+\frac{\lambda}{2}\log 2\pi\lambda+\frac{\lambda}{2}\log\psi_{2}\right]\Delta t$
$+\frac{1-\gamma}{2}\bigg\{\left[-\theta+r+(\mu-r)a(t_{k-1})-\frac{1}{2}\sigma^{2}a(t_{k-1})^{2}\right]^{2}(\Delta t)^{2}+\sigma^{2}a(t_{k-1})^{2}\left[W(t_{k})-W(t_{k-1})\right]^{2}$
$\qquad\qquad\qquad\qquad+2\left[-\theta+r+(\mu-r)a(t_{k-1})-\frac{1}{2}\sigma^{2}a(t_{k-1})^{2}\right]\sigma a(t_{k-1})\left[W(t_{k})-W(t_{k-1})\right]\Delta t\bigg\}$
$=$ $:\delta(t_{k})$
Moreover, the increments (23), (24), and (25) in the algorithm will be approximated by
$\frac{2}{T^{2}}\sum_{k=1}^{K}(T-t_{k-1})\delta(t_{k}),\ \frac{1}{T}\sum_{k=1}^{K}\frac{a(t_{k-1})-\psi_{1}}{2\psi_{2}}\delta(t_{k}),\ -\frac{1}{T}\sum_{k=1}^{K}\left(-\frac{\left(a(t_{k-1})-\psi_{1}\right)^{2}}{2}+\frac{\lambda}{2}\psi_{2}\right)\delta(t_{k}),$
respectively.
Motivated by the calculation in Stage 1, let us mainly focus on the latter two terms. In particular, recall that $a(t_{0}),\cdots,a(t_{K-1})\sim\mathcal{N}(\psi_{1},\lambda\psi_{2})$ i.i.d., hence,
$\left|\mathbb{E}\left[\frac{1}{T}\sum_{k=1}^{K}\frac{a(t_{k-1})-\psi_{1}}{2\psi_{2}}\delta(t_{k})\right]-h_{\psi,1}(\psi_{1},\psi_{2},\theta)\right|$
$=$ $\Bigg|\mathbb{E}\bigg[\frac{1}{T}\sum_{k=1}^{K}\frac{a(t_{k-1})-\psi_{1}}{2\psi_{2}}\Big[(\mu-r-\psi_{1}\sigma^{2})(a(t_{k-1})-\psi_{1})\Delta t$
$\qquad\qquad\qquad\qquad+\frac{1-\gamma}{2}\left(\cdots(\Delta t)^{2}+2\sigma^{2}\psi_{1}(a(t_{k-1})-\psi_{1})\Delta t\right)\Big]\bigg]-h_{\psi,1}(\psi_{1},\psi_{2},\theta)\Bigg|$
$=$ $\frac{|1-\gamma|}{4}\lambda\left(\mu-r-\psi_{1}\sigma^{2}\right)\left[2\left(-\theta+r+(\mu-r)\psi_{1}-\frac{1}{2}\sigma^{2}\psi_{1}^{2}\right)+3\lambda\sigma^{2}\psi_{2}\right]\Delta t,$
and
$\left|-\mathbb{E}\left[\frac{1}{T}\sum_{k=1}^{K}\left(-\frac{\left(a(t_{k-1})-\psi_{1}\right)^{2}}{2}+\frac{\lambda}{2}\psi_{2}\right)\delta(t_{k})\right]-h_{\psi,2}(\psi_{1},\psi_{2},\theta)\right|$
$=$ $\Bigg|-\mathbb{E}\bigg[\frac{1}{T}\sum_{k=1}^{K}\left(-\frac{\left(a(t_{k-1})-\psi_{1}\right)^{2}}{2}+\frac{\lambda}{2}\psi_{2}\right)\Big[-\frac{1}{2}\sigma^{2}\left(a(t_{k-1})-\psi_{1}\right)^{2}\Delta t+\frac{(a(t_{k-1})-\psi_{1})^{2}}{2\psi_{2}}\Delta t$
$\qquad\qquad\qquad\qquad+\frac{1-\gamma}{2}\left(\cdots(\Delta t)^{2}+\sigma^{2}\left(a(t_{k-1})-\psi_{1}\right)^{2}\Delta t\right)\Big]\bigg]-h_{\psi,2}(\psi_{1},\psi_{2},\theta)\Bigg|$
$=$ $\frac{|1-\gamma|}{2}\sigma^{2}\lambda^{2}\psi_{2}^{2}\Delta t.$
Moreover, we may find an upper bound for the variance of these terms. Note that across each time interval $t_{k-1},t_{k}]$, the pair $(a(t_{k-1}),\delta(t_{k}))$ are identical and independently distributed, therefore,
$\operatorname{Var}\left[\frac{1}{T}\sum_{k=1}^{K}\frac{a(t_{k-1})-\psi_{1}}{2\psi_{2}}\delta(t_{k})\right]=\frac{K}{T^{2}}\operatorname{Var}\left[\frac{a(t_{k-1})-\psi_{1}}{2\psi_{2}}\delta(t_{k})\right]$
$\leq$ $\frac{1}{4T\psi_{2}^{2}\Delta t}\mathbb{E}\left[\left(a(t_{k-1})-\psi_{1}\right)^{2}\delta(t_{k})^{2}\right]\leq\frac{\lambda\sqrt{3}}{4T\psi_{2}\Delta t}\left(\mathbb{E}\left[\delta(t_{k})^{4}\right]\right)^{1/2}.$
Similarly,
$\operatorname{Var}\left[\frac{1}{T}\sum_{k=1}^{K}\left(-\frac{\left(a(t_{k-1})-\psi_{1}\right)^{2}}{2}+\frac{\lambda}{2}\psi_{2}\right)\delta(t_{k})\right]=\frac{K}{T^{2}}\operatorname{Var}\left[\left(-\frac{\left(a(t_{k-1})-\psi_{1}\right)^{2}}{2}+\frac{\lambda}{2}\psi_{2}\right)\delta(t_{k})\right]$
$\leq$ $\frac{1}{4T\Delta t}\mathbb{E}\left[\left(\left(a(t_{k-1})-\psi_{1}\right)^{2}-\lambda\psi_{2}\right)^{2}\delta(t_{k})^{2}\right]\leq\frac{\lambda^{2}\psi_{2}^{2}\sqrt{15}}{2T\Delta t}\left(\mathbb{E}\left[\delta(t_{k})^{4}\right]\right)^{1/2}.$
By the definition of $\delta(t_{k})$, we may have
$\mathbb{E}\left[\delta(t_{k})^{4}\right]$
$\leq$ $C\mathbb{E}\Bigg[\left(1+|\theta|^{4}+a(t_{k-1})^{8}+(\lambda\log\lambda)^{4}+(\psi_{2}^{-1}+\log\psi_{2})^{4}+\psi_{1}^{8}\right)(\Delta t)^{2}+\sigma^{4}a(t_{k-1})^{4}(W(t_{k})-W(t_{k-1}))^{4}$
$\qquad\qquad+\left(1+|\theta|^{8}+a(t_{k-1})^{16}\right)(\Delta t)^{8}+\sigma^{8}a(t_{k-1})^{8}(W(t_{k})-W(t_{k-1}))^{8}$
$\qquad\qquad+(1+|\theta|^{4}+a(t_{k-1})^{8})a(t_{k-1})^{4}(\Delta t)^{4}(W(t_{k})-W(t_{k-1}))^{4}\Bigg]$
$\leq$ $C\mathbb{E}\Bigg[\left(1+|\theta|^{4}+a(t_{k-1})^{8}+(\lambda\log\lambda)^{4}+(\psi_{2}^{-1}+\log\psi_{2})^{4}+\psi_{1}^{8}\right)(\Delta t)^{2}+\left(1+|\theta|^{8}+a(t_{k-1})^{16}\right)(\Delta t)^{8}$
$\qquad\qquad+(1+|\theta|^{4}+a(t_{k-1})^{8})a(t_{k-1})^{4}(\Delta t)^{6}\Bigg]$
$\leq$ $C\left(1+|\theta|^{4}+(\lambda\log\lambda)^{4}+(\psi_{2}^{-1}+\log\psi_{2})^{4}+\psi_{1}^{8}+\lambda^{4}\psi_{2}^{4}\right)(\Delta t)^{2}$
$+C\left(1+|\theta|^{8}+\psi_{1}^{16}+\lambda^{8}\psi_{2}^{8}\right)(\Delta t)^{8}+C\left(1+|\theta|^{4}+\psi_{1}^{8}+\lambda^{4}\psi_{2}^{4}\right)\lambda^{2}\psi_{2}^{2}(\Delta t)^{6}$
where $C$ is a constant that only depends on model primitives $\mu,r,\sigma,\gamma$.
From the above calculation, the recursive relation of the sequence in the learning procedure can still be written as the same as (38), except that the “error” term $\xi_{1,i+1}$ and $\xi_{2,i+1}$ satisfy
$\beta_{1,i}$= $\left|\mathbb{E}\left[\xi_{1,i+1}|\theta_{i},\boldsymbol{\psi}_{i}\right]\right|\leq C\lambda\left(1+|\psi_{1,i}|^{3}+3\lambda\psi_{2,i}+3\lambda|\psi_{1,i}|\psi_{2,i}\right)\Delta t_{i}$
$\beta_{2,i}$= $\left|\mathbb{E}\left[\xi_{2,i+1}|\theta_{i},\boldsymbol{\psi}_{i}\right]\right|\leq C\lambda^{2}\psi_{2,i}^{2}\Delta t_{i}$
$\zeta_{1,i}$= $\operatorname{Var}\left[\xi_{1,i+1}|\theta_{i},\boldsymbol{\psi}_{i}\right]\leq C\lambda\psi_{2,i}^{-1}\sqrt{d_{1,i}+d_{2,i}(\Delta t_{i})^{4}+d_{3,i}(\Delta t_{i})^{6}}$
$\zeta_{2,i}$= $\operatorname{Var}\left[\xi_{2,i+1}|\theta_{i},\boldsymbol{\psi}_{i}\right]\leq C\lambda^{2}\psi_{2,i}^{2}\sqrt{d_{1,i}+d_{2,i}(\Delta t_{i})^{4}+d_{3,i}(\Delta t_{i})^{6}},$
where
$d_{1,i}=1+(\lambda\log\lambda)^{4}+b_{i}^{4}+c_{i}^{8}+\lambda^{4}c_{i}^{4}$
$d_{2,i}=1+c_{i}^{16}+\lambda^{8}c_{i}^{8}$
$d_{3,i}=\lambda c_{i}^{2}(1+c_{i}^{8}+\lambda c_{i}^{4}).$
Following the parallel calculation in the Stage 2, for $i\geq I$ that is sufficiently large, we have
$\leq$ $(\psi_{1,i}-\psi_{1}^{*})^{2}+2a_{\psi,i}(\psi_{1,i}-\psi_{1}^{*})\left[h_{\psi,1}(\psi_{1,i},\psi_{2,i},\theta_{i})+\beta_{1,i}\right]+a_{\psi,i}^{2}h_{\psi,1}(\psi_{1,i},\psi_{2,i},\theta_{i})^{2}+a_{\psi,i}^{2}(\beta_{1,i}^{2}+\zeta_{1,i})$
$+2a_{\psi,i}^{2}h_{\psi,1}(\psi_{1,i},\psi_{2,i},\theta_{i})\beta_{1,i}$
$=$ $(1-a_{\psi,i}\lambda\gamma\sigma^{2})(\psi_{1,i}-\psi_{1}^{*})^{2}+\frac{1}{4}a_{\psi,i}^{2}\lambda^{2}\gamma^{2}\sigma^{4}(\psi_{1,i}-\psi_{1}^{*})^{2}+a_{\psi,i}^{2}(\beta_{1,i}^{2}+\zeta_{1,i})$
$+a_{\psi,i}(1-\frac{\lambda\gamma\sigma^{2}}{2}a_{\psi,i})\left[(\psi_{1,i}-\psi_{1}^{*})^{2}+1\right]\beta_{1,i}$
$\leq$ $(1-\frac{1}{2}a_{\psi,i}\lambda\gamma\sigma^{2})(1-\frac{1}{2}a_{\psi,i}\lambda\gamma\sigma^{2}+a_{\psi,i}\beta_{1,i})(\psi_{1,i}-\psi_{1}^{*})^{2}+a_{\psi,i}(1-\frac{\lambda\gamma\sigma^{2}}{2}a_{\psi,i})\beta_{1,i}+a_{\psi,i}^{2}(\beta_{1,i}^{2}+\zeta_{1,i}).$
Without loss of generality, we may take $I$ large enough such that $1-\frac{1}{2}a_{\psi,i}\lambda\sigma^{2}+a_{\psi,i}\beta_{1,i}\in(0,1)$, and $1-\frac{1}{2}a_{\psi,i}\lambda\sigma^{2}\in(0,1)$. Moreover, given our condition, $a_{\psi,n}\sim\frac{1}{n}$, $b_{n},c_{n}\sim\log n$, and $\Delta t_{i}\sim\frac{1}{n}$, we can also ensure $d_{2,i}(\Delta t_{i})^{4}\leq d_{1,i}$, $d_{3,i}(\Delta t_{i})^{6}\leq d_{1,i}$ for all $i\geq I$. Then, we obtain
$\leq$ $(1-\frac{1}{2}a_{\psi,i}\lambda\sigma^{2})(\psi_{1,i}-\psi_{1}^{*})^{2}+a_{\psi,i}\beta_{1,i}+a_{\psi,i}^{2}(\beta_{1,i}^{2}+\zeta_{1,i})$
$\leq$ $(1-\frac{1}{2}a_{\psi,i}\lambda\sigma^{2})(\psi_{1,i}-\psi_{1}^{*})^{2}+a_{\psi,i}C\lambda\left(1+c_{i}^{3}+3\lambda c_{i}^{2}\right)\Delta t_{i}$
$+a_{\psi,i}^{2}C\left[\lambda^{2}\left(1+c_{i}^{3}+3\lambda c_{i}^{2}\right)^{2}(\Delta t_{i})^{2}+\lambda b_{i}\sqrt{d_{1,i}}\right]$
$\leq$ $(1-\frac{1}{2}a_{\psi,i}\lambda\sigma^{2})(\psi_{1,i}-\psi_{1}^{*})^{2}+a_{\psi,i}^{2}C\lambda\left[(1+c_{i}^{3}+3\lambda c_{i}^{2})\frac{\Delta t_{i}}{a_{\psi,i}}+\lambda\left(1+c_{i}^{3}+3\lambda c_{i}^{2}\right)^{2}(\Delta t_{i})^{2}+b_{i}\sqrt{d_{1,i}}\right]$
$\leq$ $(1-\frac{1}{2}a_{\psi,i}\lambda\sigma^{2})(\psi_{1,i}-\psi_{1}^{*})^{2}+a_{\psi,i}^{2}C\lambda(\log i)^{5}.$
By the similar argument in the end of Stage 2 and Lemma 4, our desired conclusion still holds.